The Mathematics Club of the Ohio Science High School has six members. How many ways can this club select a president and secretary assuming that all members are eligible but no one can hold both positions?


1
Expert's answer
2022-05-16T11:28:49-0400

If the order is important we use permutations

"^{6}P_2=\dfrac{6!}{(6-2)!}=6(5)=30"